I am thinking of a number.
I multiply it by 2 and add 104.
I get the same answer if I multiply by 10 and subtract 8.
What is my number?

Answer by amarjeeth123(569)

Let the number be x.
We have 2x+104=10x-8
10x-2x=104+8
8x=112
x=112/8=14.
The number is 14.